Can i have your help with my math.
1.Gary is x years old today. two years ago, his grandfather was 3 times older than Gary at that time.
a.) the age difference of Gary and his grandfather is 48 years. write a linear equation to represent this age difference.
b.) Find Gary's grandfather's age today.

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

1.Gary is g years old today. two years ago, his grandfather was 3 times older than Gary at that time.
Gary = g yrs
f-2 = 3(g-2)
a.) the age difference of Gary and his grandfather is 48 years. write a linear equation to represent this age difference.
3(g-2)+2 - g = 48
--------------
b.) Find Gary's grandfather's age today.
3g-6+2 - g = 48
2g - 4 = 48
g = 24+2
g = 26 yrs (Gary's age now)
---
Solve for "f":
f = 3(g-2)+2
f = 3g-4
f = 3*26-4 = 74 yrs (Grandfather's age now)
===============
Cheers,
Stan H.
===============

Answer by richwmiller(17219)

g-2=3(x-2)
a) g-x=48
g=48+x
g-48=x
b)
g-2=3((g-48)-2)
g-2=3*(g-50)
g-2=3g-150
148=2g
74=g grandfather
g-48=x
74-48=26
Gary x=26